OpenStudy (anonymous):

If z=x-iy and w=y-ix, find all values of z which satisfy z^2=w. (note:a+bi=c+di and only if a=b and c=d.)

OpenStudy (anonymous):

we have: z = x-iy w = y-ix z^2 = w thus (x-iy)^2 = y-ix x^2 - 2xiy - y^2 = y-ix obviously this will have quite a few solutions, so it's k to leave it in this form. can you solve it?

OpenStudy (anonymous):

Okay. I think I may be able to solve it. I am not sure where exactly to go from here.

OpenStudy (anonymous):

just play around with it, you might get somewhere

OpenStudy (anonymous):

Okay. I will do that. Thank you.

OpenStudy (anonymous):

\[ x^2 - 2xiy - y^2 = y-ix \implies \color{red}{x^2-y^2} + \color{blue}{-2xy}i = \color{red}y+\color{blue}{-x}i \]

OpenStudy (anonymous):

\[ -2xy = -x\\ x^2-y^2=y \]

OpenStudy (anonymous):

When \(x=0\), we get \[ 0=0\\ -y^2 = y \]When \(y=0\) then both work, so \((x,y) = (0,0)\) is one possible solution. We if \(x=0,y\neq 0\) then: \[ -y^2=y\implies -y = 1 \implies y=-1 \]So \((x,y) = (0,-1)\) is another solution. Then consider \(x\neq 0\).\[ -2xy=-x \implies-2y = -1 \implies y=\frac 12\\ x^2 - \left(\frac 12 \right)^2 = \frac{1}{2} \implies x^2 = \frac 34\implies x = \pm \frac{\sqrt 3}{2} \]

OpenStudy (anonymous):

It helps to set up two cases where a variable is zero and where a variable is not 0, so that you can divide by that variable.

OpenStudy (anonymous):

The solutions I am getting are \[ (x,y) \in \{(0,0),(0,-1), (\sqrt 3/2,1/2), (-\sqrt 3/2,1/2)\} \]

OpenStudy (anonymous):

Oh yeah, \((x,y) \to x-yi =z\).

OpenStudy (anonymous):

For example \(z=0\), \(z=i\), and \(z=(\sqrt 3 -i)/2\).

OpenStudy (anonymous):

Okay. Awesome. You have truly been a lifesaver. Thank you so much! I appreciate it! The answer would be the solution set you gave correct? Thank you so much again!

OpenStudy (anonymous):

I would double check that the solutions work

OpenStudy (anonymous):

Okay. I will. Thank you again! (: